OpenStudy (bellezabrilla):

Can someone tell me if I'm correct or not ? The question asks which equation is equivalent to 2+3(2x+4)=13 and for my equation I got 14+6x+13

OpenStudy (the1loveguru):

2+ 6x +12 =13 14 + 6x = 13 6x = -14 +13 6x = -1 x = -1/6

OpenStudy (bellezabrilla):

But the answer choices doesn't have any choices that says x=

OpenStudy (bellezabrilla):

The answer choices are A.6x+6=13 B.6x+14=13 C.10x+4=13 D.10x+20=13

OpenStudy (bellezabrilla):

I think it's B

OpenStudy (the1loveguru):

I would choose B too